OpenStudy (anonymous):

y= - cos(pi/3(x) + pi) Find the amplitude, period, and phase shift please

OpenStudy (anonymous):

amplitude is easiest, it is the absolute value of the coefficient out front since \(|-1|=1\) the amplitude is \(1\)

OpenStudy (anonymous):

period is \(\frac{2\pi}{b}\) in your case \(b=\frac{\pi}{3}\)

OpenStudy (anonymous):

compute \[\frac{2\pi}{\frac{\pi}{3}}\]to get your period

OpenStudy (anonymous):

i don't know how to find the phase shift with a formula maybe @freckles does

OpenStudy (anonymous):

you can find it without a formula by setting \[\frac{\pi}{3}x+\pi=0\] and solve for \(x\)

OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

@freckles could you help me find the phase shift?

OpenStudy (freckles):

the way satellite suggested is the way I would find phase shift

OpenStudy (anonymous):

Okay thank you!!

OpenStudy (anonymous):

Does -3 sound right?

OpenStudy (anonymous):

@freckles or @satellite73

OpenStudy (anonymous):

Nevermind